OpenStudy (anonymous):

2x^2 - 15x + 25 =0 ?

OpenStudy (amistre64):

2*25 = 50 -10*-5 = 50 -10-5 = -15 (x-10/2) (x-5/2) x = 5 , 5/2 perhaps?

OpenStudy (anonymous):

You should use quadratic formula.

OpenStudy (anonymous):

If you factor it. (x-5)(2x-5) = 0 So x=5 or 5/2